I have a word problem that I have been trying to understand:
The sum of three numbers is 150
o The largest is 25 times the smallest
o While the sum of the smallest and three times the largest is 50.
o What are my numbers?
I my problem set up like x+y+z=150

Answer by Alan3354(69443)

The sum of three numbers is 150
o The largest is 25 times the smallest
o While the sum of the smallest and three times the largest is 50.
o What are my numbers?
I my problem set up like x+y+z=150
----------------
That's true,
x + y + z = 150
x = smallest, z = largest
--> z = 25x (why didn't you mention that?)
x + 3z = 50 (or that?)
=======================
Can you do it now?